### v1

@@ Abstract
Keywords:
Relative Pose Estimation Radial Distortion RANSAC
@@ 1 Introduction
Estimating the relative pose of two cameras, i.e., estimating the relative rotation, translation, and potentially internal calibration parameters of both cameras, is a fundamental problem in computer vision. Relative pose solvers are core components of Structure-from-Motion (SfM) [46, 39] and localization pipelines [35, 44, 34] and play an important role in robotics [29, 30] and autonomous driving [38].
@@
A predominant way to estimate the relative pose of two cameras is based on 2D-2D point correspondences between the two images. Due to noise and the presence of outliers, robust estimation algorithms, such as RANdom SAmple Consensus (RANSAC) [9], or its more modern variants [33, 2], are used for the estimation. Inside RANSAC two different steps are performed: (i) Estimating the camera geometry from a (small or minimal) sample of correspondences and classifying all correspondences into inliers and outliers w.r.t. the obtained camera model. (ii) Local optimization (LO) of the camera model parameters on (a subset of) the inliers to better account for noise in the 2D point positions [6, 25].
@@
The main objective of the first step is to obtain a camera geometry estimate and the subset of correspondences consistent with it. Small samples are preferable since the number of RANSAC iterations, and thus the run-time, depends exponentially on the number of correspondences required for model estimation. Solvers that estimate the camera geometry using a minimal number of correspondences and using all available polynomial constraints are known as minimal solvers. The most commonly used minimal solvers for relative pose estimation are the well-known 5-point solver [31] for calibrated cameras and the 7-point solver [12] for uncalibrated cameras. Both are highly efficient.
@@
Minimal solvers produce estimates that perfectly fit the correspondences in the minimal sample. In practice, the 2D point correspondences are noisy and the noise in the 2D coordinates propagates to the estimates. The goal of the second step, i.e., LO inside RANSAC, is to reduce the impact of measurement noise on pose accuracy [25]. Commonly, a non-minimal solver that fits model parameters to a larger-than-minimal sample is used [6], or a robust cost function that optimizes model parameters on all inliers is minimized [25]. The most common non-minimal solver for relative pose estimation is the linear 8-point solver [12].
@@
All previously mentioned solvers, i.e., the 5-point, 7-point, and 8-point solvers, are widely used in SfM pipelines and other applications. They are based on the pinhole camera model. Yet, virtually all cameras exhibit some amount of radial distortion. Ignoring the distortion, even for standard consumer cameras, can lead to errors in 3D reconstruction [10], camera calibration accuracy, etc.
@@
There are several ways to deal with radial distortion: (1) Ignore radial distortion estimation during RANSAC and model it only in a post-processing step, e.g., during bundle adjustment in SfM [43, 40]. (2) Ignore radial distortion in the first RANSAC step but take it into account in the second step (LO), e.g., by using a non-minimal solver that estimates radial distortion parameters or by modeling distortion when minimizing a robust cost function. (3) Already estimate the radial distortion in the first RANSAC step (and refine it during LO). Approach (3) is the most principled solution as it enables taking radial distortion into account during inlier counting. Ignoring radial distortion inside the solver typically leads to identifying only the subset of the inliers that is less affected by the distortion. [^1] A point in one image maps to an epipolar curve in a radially distorted second image. Ignoring radial distortion thus means approximating this curve by a line. The approximation is only good around the center of strongly distorted images. As a result of only containing points that are only mildly affected by distortion, this inlier set often does not contain enough information to accurately estimate the distortion parameters. Thus, approaches (1) and (2), which operate on the inliers identified beforehand, are likely to fall into local minima, without recovering correct distortion and camera parameters.
@@
Radial distortion solvers are not only slower but also more complex to implement. At the same time, many of the existing minimal radial distortion solvers do not have a publicly available implementation. Even though the papers that present novel radial distortion solvers show advantages of these solvers on real data, they usually focus on presenting novel parameterizations and solution strategies and their numerical stability on synthetic data. Real experiments are mostly limited to small datasets (e.g., a single scene), simpler variants of RANSAC, and qualitative instead of quantitative results. The above-mentioned facts are most likely the reasons why (minimal) radial distortion solvers are not often used in practice. Instead, it is common to use either approach (1) or (2) [40, 43]. Naturally, this leads to the question whether (minimal) radial distortion solvers are actually necessary in practical applications.
@@ 2 Related Work
Equal and unknown radial distortion: Fitzgibbon [10] introduced an one-parameter division model for modeling distortion and an algorithm for estimating the fundamental matrix with equal and unknown radial distortion using this model. This algorithm does not use the singularity constraint on the fundamental matrix, necessitating 9 point correspondences instead of the minimal 8. This approach simplifies the problem into a standard quadratic eigenvalue problem with up to 10 solutions. The first minimal solution for epipolar geometry estimation with the one-parameter division model using 8 point correspondences was proposed by Kukelova et al. [21], using the Gröbner basis method [7] to solve a system of polynomial equations. This solver has been improved by using an automatic generator of Gröbner basis solvers [17], performing Gauss-Jordan (G-J) elimination of a 32×48 matrix and eigenvalue decomposition of a 16×16 matrix, which has up to 16 solutions. In [13], Jiang et al. used 7 point correspondences to solve the problem of essential matrix estimation for two cameras with equal and unknown focal length and radial distortion. This problem results in a complex system of polynomial equations and a large solver that performs the LU decomposition of an 886×1011 matrix and computes the eigenvalues of a 68×68 matrix. Thus, this solver is too time-consuming for practical applications.
@@
One unknown radial distortion: In [16], Kuang et al. studied three minimal cases for relative pose estimation with a single unknown radial distortion based on the Gröbner basis method: 8-point fundamental matrix and radial distortion; 7-point essential matrix, focal length and radial distortion; 6-point essential matrix and radial distortion. However, these solvers assume one of the two cameras has known or no radial distortion, which may not be realistic in practice.
[…]
@@
Recently, Oskarsson [32] presented a unified formulation for relative pose problems involving radial distortion and proposed more efficient minimal solvers for all different configurations. While some of the proposed solvers are already quite efficient, e.g., the 8-point solver for uncalibrated cameras with common radial distortion, others, like the 9-point solver for different distortions, are still too slow and/or numerically unstable to be useful in practice.
@@
Parameter sampling: Instead of jointly estimating the absolute camera pose and the focal length of an uncalibrated camera, Sattler et al. [36] proposed a RANSAC variant that combines parameter sampling and parameter estimation. In each RANSAC iteration, they first randomly sample a focal length value and then estimate the pose of the now-calibrated camera. The probability distribution over the focal length values is then updated based on the number of inliers of the estimated pose. We propose a simpler sampling-based strategy for relative pose estimation that uses a small fixed set of undistortion parameters. In contrast to [36], our approach can easily be applied to 2D sampling problems, e.g., two different and unknown distortion parameters.
@@ 3 Radial Distortion Solvers
$u(\mathbf{x}_{i},\mathbf{\Lambda})^{\top}\mathbf{F}u(\mathbf{x}_{i}^{\prime},\mathbf{\Lambda}^{\prime})=0\enspace,$ (1)
@@
In this paper, we model the undistortion function using the one-parameter division model [10]. In this model, given an observed radially distorted point with homogeneous coordinates $\mathbf{x}=[x_{d},y_{d},1]^{\top}$, and the undistortion parameter $\lambda\in\mathbb{R}$, the undistorted image point is given as
$u(\mathbf{x},\lambda)=[x_{d},y_{d},1+\lambda(x_{d}^{2}+y_{d}^{2})]^{\top}\enspace,$ (2)
@@
  • 8pt $\mathbf{F}\lambda$ : Among all minimal 8pt solvers [21, 17, 24, 32], the solver from [32] is the most efficient. It formulates the elements of $\mathbf{F}$ as functions of the distortion parameter $\lambda$, and obtains an univariate polynomial in $\lambda$ of degree 16, which can be solved using the Sturm sequences, with up to 16 solutions.
  • 9pt $\mathbf{F}\lambda$ : By ignoring the $\det(\mathbf{F})=0$ constraint, the $\mathbf{F}\lambda$ problem can be solved using nine point correspondences. [10] solves the nine equations (1) by converting them into a polynomial eigenvalue problem. While [10] was able to remove several spurious solutions by transforming the original eigenvalue problem of size $18\times 18$ into a problem of size $10\times 10$[10] also observed that 4 of the 10 solutions of this system are imaginary. In this paper, we propose a modification of the solver proposed in [10], in which we directly remove 4 imaginary solutions, resulting in a more efficient solver that needs to find the eigenvalues of a smaller $6\times 6$ matrix. To remove these 4 imaginary solutions, we use the structure of matrices that appear in the polynomial eigenvalue formulation of this problem and the method proposed in [18]. For more details, see the Supplementary Material (SM).
@@
  • 9pt $\mathbf{F}\lambda_{1}\lambda_{2}$ : Equations for cameras with different unknown distortions are more complex than for the equal distortion case. In this case the system of equations has 24 solutions and the fastest Gröbner basis solver from [24], which returns 24 solutions, performs elimination of a large matrix of size $84\times 117$ followed by the eigendecomposition of a $24\times 24$ matrix. The recently published parameterization of this problem in [32] performs elimination of a smaller matrix of size $51\times 99$ followed by the eigendecomposition of a $48\times 48$ matrix. The solver returns up to 48 solutions. However, it is still faster than the solver from [24]. Thus, in our experiments, we use the solver from [32].
  • 10pt $\mathbf{F}\lambda_{1}\lambda_{2}$ : In [20] it was shown that in many scenarios inside RANSAC it is preferable to sample 10 instead of 9 points and run the more efficient 10pt solver. In [20], the authors proposed several variants of the 10pt solver. In this paper, we use the variant based on a Gröbner basis, made available by the authors. The 10pt solvers cannot be easily modified to work with more than 10 points, and thus we use this solver only in the first step of RANSAC, i.e., instead of the minimal solver, and not in the LO step.
  • 12pt $\mathbf{F}\lambda_{1}\lambda_{2}$ : Similarly to the 9pt $\mathbf{F}\lambda$ solver, we can ignore the $\det(\mathbf{F})=0$ constraint in the case of different radial distortions and formulate the problem as a polynomial eigenvalue problem. Such a solver using 12 point correspondences was proposed in [19]. The solver performs the eigendecomposition of a $12\times 12$ matrix. Yet, as observed in [19], it has eight infinite eigenvalues, and at most four finite real solutions. In this paper, we modify this solver. We remove all infinite eigenvalues in advance. Thus, in our new variant of the solver, we only need to find the eigenvalues of a $4\times 4$ matrix. For more details, see the SM. This solver, as well as the 9pt $\mathbf{F}\lambda$ solver, can be easily extended to more points and used in the LO step of RANSAC.
@@ 4 Experiments
Datasets. We evaluate the different approaches for radial distortion relative pose estimation on three datasets: Phototourism [15], ETH3D [41], and our new benchmark, each covering different scenarios. Phototourism contains images of popular landmarks downloaded from Internet photo collections, taken with a wide variety of cameras. Phototourism provides ground truth camera poses obtained via SfM and undistorted images with known intrinsics. The dataset is widely used for evaluating local features and feature matching [15]. For our experiments, we used 13 scenes with 5,000 image pairs per scene. It covers indoor and outdoor scenes captured with a DSLR camera and a multi-camera rig containing synchronized cameras with varying field-of-views. Unlike Phototourism, which contains images taken at different points in time and under varying lighting conditions, the images were taken in a short period of time for each scene. Ground truth poses were obtained by aligning the images to high-precision laser scans. ETH3D provides undistorted images together with their intrinsic calibration. We use 2,037 image pairs from 12 ETH3D scenes.
@@
Existing datasets [42, 41] containing radially distorted images mostly involve only one or two different camera lenses with little variation in the distortion parameters. Testing the sampling-based strategy on such images could be biased, as it would have been as good as the distance of the used sampled value from the one/two ground truth values. We thus created a new benchmark with a higher variation in the distortion parameters, consisting of two scenes: ROTUNDA and CATHERAL. For both scenes, we build upon previously recorded images [20, 37], taken by GoPro cameras and kindly provided by the authors. We added images from phone cameras with very small lens distortion, images taken by action cameras (GoPro, Insta360 Ace Pro) with different distortions, and a few images downloaded from Flickr. We created ground truth for the benchmark (poses and the parameters of the one-parameter division model) using the RealityCapture SfM software [1]. We use 3,424 image pairs with $\lambda_{1}\neq\lambda_{2}$ and 1,795 image pairs with $\lambda_{1}=\lambda_{2}$ for ROTUNDA and 10k sampled image pairs for both $\lambda_{1}\neq\lambda_{2}$ and $\lambda_{1}=\lambda_{2}$ for CATHERAL. The scenes are visualized in the SM.
@@
Evaluation measures. Following [15], given the ground truth and the estimated relative pose, we measure the rotation error and the translation error. The rotation error is defined as the angle of the rotation matrix aligning the estimated with the ground truth rotation matrix. The translation error is defined as the angle between the estimated and the ground truth translation vector. Finally, the pose error is defined as the maximum of the rotation error and the translation error. To obtain poses from the estimated fundamental matrices we use ground truth intrinsics [15]. We report the average (AVG) and median (MED) pose errors, as well as the area under the recall curve (AUC) for different thresholds on the pose error. In addition, we measure the distortion error $\epsilon(\lambda)$ as the absolute distance between ground truth and estimated distortion parameters. For the $\mathbf{F}\lambda_{1}\lambda_{2}$ problem, we measure the distortion error as $0.5\cdot(\epsilon(\lambda_{1})+\epsilon(\lambda_{2}))$.
@@
Experimental setup. We obtain point correspondences between images by matching Superpoint [8] features with the LightGlue (LG) [26] matcher. Experiments with SIFT [28] and LG can be found in the SM.
@@
In the first step of RANSAC, we use the minimal 7pt $\mathbf{F}$, 8pt $\mathbf{F}\lambda$, 9pt $\mathbf{F}\lambda_{1}\lambda_{2}$, and the non-minmal 9pt $\mathbf{F}\lambda$, and 10pt $\mathbf{F}\lambda_{1}\lambda_{2}$ solvers (cf. Sec. 3), as well as the sampling strategies that combine the 7pt solver with a set of pre-defined undistortion parameters (cf. Sec. 3.1). We denote the latter by appending the list of parameters, e.g., $\{0,-0.6,-1.2\}$, to the solver configuration. We integrate the solvers and strategies into two modern RANSAC implementations: Graph-Cut RANSAC (GC-RANSAC) [2], and Poselib [23]. We chose these two variants as they use different approaches for Local Optimization (LO). GC-RANSAC uses a non-minimal solver for fitting the relative pose and distortion parameter(s) to a larger-than-minimal sample when doing final pose polishing on all inliers and in the local optimization step. We use the refined 9-point solver (9pt $\mathbf{F}\lambda$) for the equal and unknown radial distortion case ($\lambda_{1}\,=\,\lambda_{2}$), and the refined 12-point solver (12pt $\mathbf{F}\lambda_{1}\lambda_{2}$) for the different and unknown radial distortion case ($\lambda_{1}\,\not=\,\lambda_{2}$). In contrast, the LO step in PoseLib relies on Levenberg-Marquardt (LM) optimization of the truncated Tangent Sampson Error [45], starting from the estimate provided by the minimal solver / sampling strategy. The pose returned by RANSAC is further polished by LM optimization over all inliers. For Poselib we also include experiments with a version corresponding to Approach (1) mentioned in Sec. 1, i.e., modelling the radial distortion only in the final optimization. We denote this approach as 7pt $\mathbf{F}$ with $\mathbf{F}$+ $\mathbf{F}\lambda$ and $\mathbf{F}$+ $\mathbf{F}\lambda_{1}\lambda_{2}$ for refinement. To determine which points are inliers we use the Tangent Sampson Error [45] with a fixed threshold (1 px in GC-RANSAC, 3 px in Poselib). Using the Tangent Sampson Error is important since the standard Sampson Error in undistorted images leads to a radial bias in the optimization [45].
@@
We use normalized image coordinates from the range $[-0.5,0.5]^{2}$, obtained by subtracting the image center and dividing by the length of the longer image side. For this normalization, the undistortion parameter should be greater than $-2$, as otherwise the distortion would mirror the image. In RANSAC, we discard models with radial distortion outside of the plausible range $\left[-2.0,0.5\right]$.

### v2

@@ 1 Introduction
Estimating the relative pose of two cameras, i.e., estimating the relative rotation, translation, and potentially internal calibration parameters of both cameras, is a fundamental problem in computer vision. Relative pose solvers are core components of Structure-from-Motion (SfM) [wu2011visualsfm, Schoenberger2016CVPR] and localization pipelines [sattler2016efficient, svarm2016city, sarlin2019coarse] and play an important role in robotics [mur2015orb, mur2017orb] and autonomous driving [scaramuzza2011visual].
@@
A predominant way to estimate the relative pose of two cameras is based on 2D-2D point correspondences between the two images. Due to noise and the presence of outliers, robust estimation algorithms, such as RANdom SAmple Consensus (RANSAC) [Fischler-Bolles-ACM-1981], or its more modern variants [raguram2013usac, barath2017graph], are used for the estimation. Inside RANSAC two different steps are performed: (i) Estimating the camera geometry from a (small or minimal) sample of correspondences and classifying all correspondences into inliers and outliers w.r.t. the obtained camera model. (ii) Local optimization (LO) of the camera model parameters on (a subset of) the inliers to better account for noise in the 2D point positions [Chum-2003, lebeda2012fixing].
@@
The main objective of the first step is to obtain a camera geometry estimate and the subset of correspondences consistent with it. Small samples are preferable since the number of RANSAC iterations, and thus the run-time, depends exponentially on the number of correspondences required for model estimation. Solvers that estimate the camera geometry using a minimal number of correspondences and using all available polynomial constraints are known as minimal solvers. The most commonly used minimal solvers for relative pose estimation are the well-known 5-point solver [nister2004efficient] for calibrated cameras and the 7-point solver [hartley2003multiple] for uncalibrated cameras. Both are highly efficient.
@@
Minimal solvers produce estimates that perfectly fit the correspondences in the minimal sample. In practice, the 2D point correspondences are noisy and the noise in the 2D coordinates propagates to the estimates. The goal of the second step, i.e., LO inside RANSAC, is to reduce the impact of measurement noise on pose accuracy [lebeda2012fixing]. Commonly, a non-minimal solver that fits model parameters to a larger-than-minimal sample is used [Chum-2003], or a robust cost function that optimizes model parameters on all inliers is minimized [lebeda2012fixing]. The most common non-minimal solver for relative pose estimation is the linear 8-point solver [hartley2003multiple].
@@
All previously mentioned solvers, i.e., the 5-point, 7-point, and 8-point solvers, are widely used in SfM pipelines and other applications. They are based on the pinhole camera model. Yet, virtually all cameras exhibit some amount of radial distortion. Ignoring the distortion, even for standard consumer cameras, can lead to errors in 3D reconstruction [fitzgibbon2001simultaneous], camera calibration accuracy, etc.
@@
There are several ways to deal with radial distortion: (1) Ignore radial distortion estimation during RANSAC and model it only in a post-processing step, e.g., during bundle adjustment in SfM [snavely2008modeling, schonberger2016structure]. (2) Ignore radial distortion in the first RANSAC step but take it into account in the second step (LO), e.g., by using a non-minimal solver that estimates radial distortion parameters or by modeling distortion when minimizing a robust cost function. (3) Already estimate the radial distortion in the first RANSAC step (and refine it during LO). Approach (3) is the most principled solution as it enables taking radial distortion into account during inlier counting. Ignoring radial distortion inside the solver typically leads to identifying only the subset of the inliers that is less affected by the distortion. [^1] A point in one image maps to an epipolar curve in a radially distorted second image. Ignoring radial distortion thus means approximating this curve by a line. The approximation is only good around the center of strongly distorted images. As a result of only containing points that are only mildly affected by distortion, this inlier set often does not contain enough information to accurately estimate the distortion parameters. Thus, approaches (1) and (2), which operate on the inliers identified beforehand, are likely to fall into local minima, without recovering correct distortion and camera parameters.
@@
Radial distortion solvers are not only slower but also more complex to implement. At the same time, many of the existing minimal radial distortion solvers do not have a publicly available implementation. Even though the papers that present novel radial distortion solvers show advantages of these solvers on real data, they usually focus on presenting novel parameterizations and solution strategies and their numerical stability on synthetic data. Real experiments are mostly limited to small datasets (e.g., a single scene), simpler variants of RANSAC, and qualitative instead of quantitative results. The above-mentioned facts are most likely the reasons why (minimal) radial distortion solvers are not often used in practice. Instead, it is common to use either approach (1) or (2) [schonberger2016structure, snavely2008modeling]. Naturally, this leads to the question whether (minimal) radial distortion solvers are actually necessary in practical applications.
@@ 2 Related Work
Equal and unknown radial distortion: Fitzgibbon [fitzgibbon2001simultaneous] introduced an one-parameter division model for modeling distortion and an algorithm for estimating the fundamental matrix with equal and unknown radial distortion using this model. This algorithm does not use the singularity constraint on the fundamental matrix, necessitating 9 point correspondences instead of the minimal 8. This approach simplifies the problem into a standard quadratic eigenvalue problem with up to 10 solutions. The first minimal solution for epipolar geometry estimation with the one-parameter division model using 8 point correspondences was proposed by Kukelova et al[kukelova2007minimal], using the Gröbner basis method [cox2005using] to solve a system of polynomial equations. This solver has been improved by using an automatic generator of Gröbner basis solvers [kukelova2008automatic], performing Gauss-Jordan (G-J) elimination of a 32×48 matrix and eigenvalue decomposition of a 16×16 matrix, which has up to 16 solutions. In [jiang2015minimal], Jiang et al. used 7 point correspondences to solve the problem of essential matrix estimation for two cameras with equal and unknown focal length and radial distortion. This problem results in a complex system of polynomial equations and a large solver that performs the LU decomposition of an 886×1011 matrix and computes the eigenvalues of a 68×68 matrix. Thus, this solver is too time-consuming for practical applications.
@@
One unknown radial distortion: In [kuang2014minimal], Kuang et al. studied three minimal cases for relative pose estimation with a single unknown radial distortion based on the Gröbner basis method: 8-point fundamental matrix and radial distortion; 7-point essential matrix, focal length and radial distortion; 6-point essential matrix and radial distortion. However, these solvers assume one of the two cameras has known or no radial distortion, which may not be realistic in practice.
Different and unknown radial distortions: All of the above mentioned algorithms estimate only one radial distortion parameter for one or both cameras. In practice, e.g., using images downloaded from the Internet, two cameras can have different and unknown radial distortions. The problem of fundamental matrix estimation with different and unknown radial distortions, $\mathbf{F}\lambda_{1}\lambda_{2}$, was first studied by Barreto et al[barreto2005fundamental], proposing a non-minimal linear algorithm using 15 point correspondences (F15). The minimal 9-point case (F9) for this problem was studied in [kukelova2007two, byrod2008fast, kukelova2008automatic, kukelova2010fast]. The solver from [byrod2008fast] (F9) performs LU decomposition of a 393×389 matrix, SVD decomposition of a 69×69 matrix, and eigenvalue computation of a 24×24 matrix. In [kukelova2008automatic], a faster version based on a Gröbner basis ($\rm F9_{\rm A}$) was proposed. It performs G-J elimination of a 179×203 matrix and eigenvalue decomposition of a 24×24 matrix. However, this solver is slightly less stable than F9, and still too slow for real-time applications. Kukelova et al[kukelova2010fast] suggested an efficient, non-minimal solver using 12 point correspondences (F12) that generates up to four real solutions. However, this algorithm is more sensitive to noise than the minimal $\rm F9_{\rm A}$ [kukelova2010fast]. Balancing efficiency and noise sensitivity, Kukelova et al[kukelova2015efficient] proposed a 10-point solver that is much faster than the minimal 9-point solver and more robust to image noise than the 12-point solver.
@@
Recently, Oskarsson [Oskarsson_2021_CVPR] presented a unified formulation for relative pose problems involving radial distortion and proposed more efficient minimal solvers for all different configurations. While some of the proposed solvers are already quite efficient, e.g., the 8-point solver for uncalibrated cameras with common radial distortion, others, like the 9-point solver for different distortions, are still too slow and/or numerically unstable to be useful in practice.
@@
Parameter sampling: Instead of jointly estimating the absolute camera pose and the focal length of an uncalibrated camera, Sattler et al[Sattler2014ECCV] proposed a RANSAC variant that combines parameter sampling and parameter estimation. In each RANSAC iteration, they first randomly sample a focal length value and then estimate the pose of the now-calibrated camera. The probability distribution over the focal length values is then updated based on the number of inliers of the estimated pose. We propose a simpler sampling-based strategy for relative pose estimation that uses a small fixed set of undistortion parameters. In contrast to [Sattler2014ECCV], our approach can easily be applied to 2D sampling problems, e.g., two different and unknown distortion parameters.
@@ 3 Radial Distortion Solvers
$u(\mathbf{x}_{i},\mathbf{\Lambda})^{\top}\mathbf{F}u(\mathbf{x}_{i}^{\prime},\mathbf{\Lambda}^{\prime})=0\kern 5.0pt,$ (1)
@@
In this paper, we model the undistortion function using the one-parameter division model [fitzgibbon2001simultaneous]. In this model, given an observed radially distorted point with homogeneous coordinates $\mathbf{x}=[x_{d},y_{d},1]^{\top}$, and the undistortion parameter $\lambda\in\mathbb{R}$, the undistorted image point is given as
$u(\mathbf{x},\lambda)=[x_{d},y_{d},1+\lambda(x_{d}^{2}+y_{d}^{2})]^{\top}\kern 5.0pt,$ (2)
@@
  • 8pt $\mathbf{F}\lambda$ : Among all minimal 8pt solvers [kukelova2007minimal, kukelova2008automatic, larsson2017efficient, Oskarsson_2021_CVPR], the solver from [Oskarsson_2021_CVPR] is the most efficient. It formulates the elements of $\mathbf{F}$ as functions of the distortion parameter $\lambda$, and obtains an univariate polynomial in $\lambda$ of degree 16, which can be solved using the Sturm sequences, with up to 16 solutions.
  • 9pt $\mathbf{F}\lambda$ : By ignoring the $\det(\mathbf{F})=0$ constraint, the $\mathbf{F}\lambda$ problem can be solved using nine point correspondences. [fitzgibbon2001simultaneous] solves the nine equations (1) by converting them into a polynomial eigenvalue problem. While [fitzgibbon2001simultaneous] was able to remove several spurious solutions by transforming the original eigenvalue problem of size $18\times 18$ into a problem of size $10\times 10$[fitzgibbon2001simultaneous] also observed that 4 of the 10 solutions of this system are imaginary. In this paper, we propose a modification of the solver proposed in [fitzgibbon2001simultaneous], in which we directly remove 4 imaginary solutions, resulting in a more efficient solver that needs to find the eigenvalues of a smaller $6\times 6$ matrix. To remove these 4 imaginary solutions, we use the structure of matrices that appear in the polynomial eigenvalue formulation of this problem and the method proposed in [kukelova2012polynomial]. For more details, see the Supplementary Material (SM).
@@
  • 9pt $\mathbf{F}\lambda_{1}\lambda_{2}$ : Equations for cameras with different unknown distortions are more complex than for the equal distortion case. In this case the system of equations has 24 solutions and the fastest Gröbner basis solver from [larsson2017efficient], which returns 24 solutions, performs elimination of a large matrix of size $84\times 117$ followed by the eigendecomposition of a $24\times 24$ matrix. The recently published parameterization of this problem in [Oskarsson_2021_CVPR] performs elimination of a smaller matrix of size $51\times 99$ followed by the eigendecomposition of a $48\times 48$ matrix. The solver returns up to 48 solutions. However, it is still faster than the solver from [larsson2017efficient]. Thus, in our experiments, we use the solver from [Oskarsson_2021_CVPR].
  • 10pt $\mathbf{F}\lambda_{1}\lambda_{2}$ : In [kukelova2015efficient] it was shown that in many scenarios inside RANSAC it is preferable to sample 10 instead of 9 points and run the more efficient 10pt solver. In [kukelova2015efficient], the authors proposed several variants of the 10pt solver. In this paper, we use the variant based on a Gröbner basis, made available by the authors. The 10pt solvers cannot be easily modified to work with more than 10 points, and thus we use this solver only in the first step of RANSAC, i.e., instead of the minimal solver, and not in the LO step.
  • 12pt $\mathbf{F}\lambda_{1}\lambda_{2}$ : Similarly to the 9pt $\mathbf{F}\lambda$ solver, we can ignore the $\det(\mathbf{F})=0$ constraint in the case of different radial distortions and formulate the problem as a polynomial eigenvalue problem. Such a solver using 12 point correspondences was proposed in [kukelova2010fast]. The solver performs the eigendecomposition of a $12\times 12$ matrix. Yet, as observed in [kukelova2010fast], it has eight infinite eigenvalues, and at most four finite real solutions. In this paper, we modify this solver. We remove all infinite eigenvalues in advance. Thus, in our new variant of the solver, we only need to find the eigenvalues of a $4\times 4$ matrix. For more details, see the SM. This solver, as well as the 9pt $\mathbf{F}\lambda$ solver, can be easily extended to more points and used in the LO step of RANSAC.
@@ 4 Experiments
Datasets. We evaluate the different approaches for radial distortion relative pose estimation on three datasets: Phototourism [IMC2020], ETH3D [Schops_2017_CVPR], and our new benchmark, each covering different scenarios. Phototourism contains images of popular landmarks downloaded from Internet photo collections, taken with a wide variety of cameras. Phototourism provides ground truth camera poses obtained via SfM and undistorted images with known intrinsics. The dataset is widely used for evaluating local features and feature matching [IMC2020]. For our experiments, we used 13 scenes with 5,000 image pairs per scene. It covers indoor and outdoor scenes captured with a DSLR camera and a multi-camera rig containing synchronized cameras with varying field-of-views. Unlike Phototourism, which contains images taken at different points in time and under varying lighting conditions, the images were taken in a short period of time for each scene. Ground truth poses were obtained by aligning the images to high-precision laser scans. ETH3D provides undistorted images together with their intrinsic calibration. We use 2,037 image pairs from 12 ETH3D scenes.
@@
Existing datasets [schubert2018vidataset, Schops_2017_CVPR] containing radially distorted images mostly involve only one or two different camera lenses with little variation in the distortion parameters. Testing the sampling-based strategy on such images could be biased, as it would have been as good as the distance of the used sampled value from the one/two ground truth values. We thus created a new benchmark with a higher variation in the distortion parameters, consisting of two scenes: ROTUNDA and CATHERAL. For both scenes, we build upon previously recorded images [kukelova2015efficient, Sattler_2019_CVPR], taken by GoPro cameras and kindly provided by the authors. We added images from phone cameras with very small lens distortion, images taken by action cameras (GoPro, Insta360 Ace Pro) with different distortions, and a few images downloaded from Flickr. We created ground truth for the benchmark (poses and the parameters of the one-parameter division model) using the RealityCapture SfM software [RealityCapture]. We use 3,424 image pairs with $\lambda_{1}\neq\lambda_{2}$ and 1,795 image pairs with $\lambda_{1}=\lambda_{2}$ for ROTUNDA and 10k sampled image pairs for both $\lambda_{1}\neq\lambda_{2}$ and $\lambda_{1}=\lambda_{2}$ for CATHERAL. The scenes are visualized in the SM.
@@
Evaluation measures. Following [IMC2020], given the ground truth and the estimated relative pose, we measure the rotation error and the translation error. The rotation error is defined as the angle of the rotation matrix aligning the estimated with the ground truth rotation matrix. The translation error is defined as the angle between the estimated and the ground truth translation vector. Finally, the pose error is defined as the maximum of the rotation error and the translation error. To obtain poses from the estimated fundamental matrices we use ground truth intrinsics [IMC2020]. We report the average (AVG) and median (MED) pose errors, as well as the area under the recall curve (AUC) for different thresholds on the pose error. In addition, we measure the distortion error $\epsilon(\lambda)$ as the absolute distance between ground truth and estimated distortion parameters. For the $\mathbf{F}\lambda_{1}\lambda_{2}$ problem, we measure the distortion error as $0.5\cdot(\epsilon(\lambda_{1})+\epsilon(\lambda_{2}))$.
@@
Experimental setup. We obtain point correspondences between images by matching Superpoint [detone2018superpoint] features with the LightGlue (LG) [lindenberger2023lightglue] matcher. Experiments with SIFT [lowe2004distinctive] and LG can be found in the SM.
@@
In the first step of RANSAC, we use the minimal 7pt $\mathbf{F}$, 8pt $\mathbf{F}\lambda$, 9pt $\mathbf{F}\lambda_{1}\lambda_{2}$, and the non-minmal 9pt $\mathbf{F}\lambda$, and 10pt $\mathbf{F}\lambda_{1}\lambda_{2}$ solvers (cf. Sec. 3), as well as the sampling strategies that combine the 7pt solver with a set of pre-defined undistortion parameters (cf. Sec. 3.1). We denote the latter by appending the list of parameters, e.g., $\{0,-0.6,-1.2\}$, to the solver configuration. We integrate the solvers and strategies into two modern RANSAC implementations: Graph-Cut RANSAC (GC-RANSAC) [barath2017graph], and Poselib [PoseLib]. We chose these two variants as they use different approaches for Local Optimization (LO). GC-RANSAC uses a non-minimal solver for fitting the relative pose and distortion parameter(s) to a larger-than-minimal sample when doing final pose polishing on all inliers and in the local optimization step. We use the refined 9-point solver (9pt $\mathbf{F}\lambda$) for the equal and unknown radial distortion case ($\lambda_{1}\,=\,\lambda_{2}$), and the refined 12-point solver (12pt $\mathbf{F}\lambda_{1}\lambda_{2}$) for the different and unknown radial distortion case ($\lambda_{1}\,\not=\,\lambda_{2}$). In contrast, the LO step in PoseLib relies on Levenberg-Marquardt (LM) optimization of the truncated Tangent Sampson Error [Terekhov_2023_ICCV], starting from the estimate provided by the minimal solver / sampling strategy. The pose returned by RANSAC is further polished by LM optimization over all inliers. For Poselib we also include experiments with a version corresponding to Approach (1) mentioned in Sec. 1, i.e., modelling the radial distortion only in the final optimization. We denote this approach as 7pt $\mathbf{F}$ with $\mathbf{F}$+ $\mathbf{F}\lambda$ and $\mathbf{F}$+ $\mathbf{F}\lambda_{1}\lambda_{2}$ for refinement. To determine which points are inliers we use the Tangent Sampson Error [Terekhov_2023_ICCV] with a fixed threshold (1 px in GC-RANSAC, 3 px in Poselib). Using the Tangent Sampson Error is important since the standard Sampson Error in undistorted images leads to a radial bias in the optimization [Terekhov_2023_ICCV].
@@
We use normalized image coordinates from the range $0.5,0.5]^{2}$, obtained by subtracting the image center and dividing by the length of the longer image side. For this normalization, the undistortion parameter should be greater than 2-2, as otherwise the distortion would mirror the image. In RANSAC, we discard models with radial distortion outside of the plausible range $\left[-2.0,0.5\right]$.